# PhysLab: A Benchmark Dataset for Multi-Granularity Visual Parsing of Physics Experiments

### v1

@@ Abstract.
Visual parsing of images and videos is critical for a wide range of real-world applications. However, progress in this field is constrained by limitations of existing datasets: (1) insufficient annotation granularity, which impedes fine-grained scene understanding and high-level reasoning; (2) limited coverage of domains, particularly a lack of datasets tailored for educational scenarios; and (3) lack of explicit procedural guidance, with minimal logical rules and insufficient representation of structured task process. To address these gaps, we introduce PhysLab, the first video dataset that captures students conducting complex physics experiments. The dataset includes four representative experiments that feature diverse scientific instruments and rich human-object interaction (HOI) patterns. PhysLab comprises 620 long-form videos and provides multilevel annotations that support a variety of vision tasks, including action recognition, object detection, HOI analysis, etc. We establish strong baselines and perform extensive evaluations to highlight key challenges in the parsing of procedural educational videos. We expect PhysLab to serve as a valuable resource for advancing fine-grained visual parsing, facilitating intelligent classroom systems, and fostering closer integration between computer vision and educational technologies. The dataset and the evaluation toolkit are publicly available at https://github.com/ZMH-SDUST/PhysLab.
Procedural Video, Physics Lab Education, Visual Parsing, Multi-Granularity Annotation
Figure 1. An overview of the content of our proposed PhysLab dataset. PhysLab dataset focuses on experimental tasks conducted in complex physics laboratory settings. 1) The top row shows the data collection process and representative data samples. 2) The second row shows the distinctive operational instructions associated with physics experiments, including detailed procedures, instrument usage, and process models in Petri Net format (lower-left), along with multi-granularity annotations (lower-right).
@@ 1. Introduction
Visual parsing is a fundamental research in computer vision, encompassing a wide range of tasks such as action recognition, object detection, and HOI detection (Yu et al., 2022). It plays a pivotal role in real-world applications including surveillance, autonomous driving, and industrial inspection. In recent years, the field has achieved remarkable progress, driven by advances in deep learning and the availability of large-scale annotated datasets (Zhou et al., 2025a). However, this data-centric paradigm also exposes a strong reliance on high-quality, diverse annotations. Although existing datasets are large in scale, they remain fall short in supporting multi-task collaborative modeling and rich semantic understanding (Li et al., 2024). Specifically, they are often tailored to single-task learning or provide only coarse-grained labels, lacking the structured, fine-grained annotations needed to jointly model actions, objects, and interactions (Kim et al., 2025). Furthermore, most datasets are predominantly built around daily scenarios such as households (Tang et al., 2019), streets (Liu et al., 2024), and kitchens (Lee et al., 2024), while neglecting cognitively demanding domains like education. As a result, many baseline models tend to prioritize static appearance cues or short-term visual patterns, falling short in modeling the causal relationships, temporal dependencies, and physical constraints that are intrinsic to complex procedural tasks (Wu et al., 2024). These limitations hinder the development of models capable of deeper, human-like understanding.
@@
In contrast to conventional visual scenes, physics experiments exhibit significantly greater spatiotemporal complexity and cognitive demands (Latif et al., 2024). Experimental procedures typically involve multi-step operations, instrument configurations, and parameter adjustments, all governed by strict task dependencies and procedural order (Zou et al., 2024). These tasks also feature frequent and detailed human-instrument interactions, including tool switching and feedback observation, which require fine-grained recognition and cross-modal reasoning (Huang et al., 2024). Additionally, such processes are often underpinned by implicit physical knowledge and causal dynamics (Pols and Dekkers, 2024) that cannot be effectively captured by low-level visual features alone. To facilitate research on such complex and cognitively rich scenarios, we introduce PhysLab, the first multi-task video dataset specifically designed for physics experiment instruction.
@@
As shown in Figure 1, PhysLab focuses on four representative undergraduate physics experiments, covering a diverse set of scientific instruments and typical student behaviors. The dataset contains 620 long-form videos (totalling 31 hours), each documenting authentic student experiment sessions. Videos are segmented into action clips, and richly annotated with hierarchical labels including action categories, object locations, interaction relationships, and operation sequences. These annotations provide a unified platform for multi-granularity and multi-task learning. In addition, the dataset incorporates structured metadata encoding experimental types, procedural steps, and tools used—facilitating higher-level tasks such as procedural modeling and experimental state reasoning. In this context, a process model refers to a structured representation of the experimental workflow, capturing the logical and temporal constraints among operations. It helps in understanding how experiments unfold over time, supporting reasoning about progress, correctness, and deviations during task execution.
@@
To systematically assess the challenges introduced by PhysLab, we conducted extensive benchmark experiments spanning 8 mainstream visual parsing tasks. The evaluation results highlight the dataset’s complexity and expose substantial limitations in current methods, particularly in terms of insufficient visual context fusion, lack of prior knowledge, and pronounced inter-class performance disparity. We believe PhysLab offers a valuable foundation for advancing research in complex behavior modeling, improving educational process analysis, and bridging the gap between visual understanding and intelligent educational systems.
@@ 2.1. Temporal Parsing Datasets
Temporal parsing datasets focus on modeling the temporal dynamics of videos and support tasks such as action classification, temporal action localization, and action segmentation (Sun et al., 2022; Wang et al., 2023). Early datasets such as UCF101 (Soomro et al., 2012), HMDB51 (Kuehne et al., 2011), and Kinetics (Carreira and Zisserman, 2017) contain a variety of short, edited clips depicting everyday human actions, and have been widely adopted for action classification tasks. However, these datasets offer limited diversity in action types and lack fine-grained temporal structures. To address this, subsequent datasets like THUMOS14 (Idrees et al., 2017), ActivityNet1.3 (Caba Heilbron et al., 2015), and FineAction (Liu et al., 2022) introduced untrimmed videos with denser temporal annotations, enabling research on action detection and segmentation.
More recently, as the community shifts towards modeling structured activities, several procedural video datasets have emerged. Datasets such as Breakfast (Kuehne et al., 2014), COIN (Tang et al., 2019), and Assembly101 (Sener et al., 2022) focus on multi-step instructional tasks with well-defined goals, supporting the study of procedural understanding. In addition, datasets like EgoPER (Lee et al., 2024), CaptainCook4D (Peddi et al., 2024), and IndustReal (Schoonbeek et al., 2024) incorporate annotations of execution errors and deviations, facilitating research on anomaly detection in task workflows. Despite recent progress, existing datasets remain largely confined to narrow domains such as cooking and mechanical assembly (Tan et al., 2021). There is a significant lack of resources tailored to student laboratory experiments, which are essential for advancing research at the intersection of visual perception and intelligent educational systems (Piloto et al., 2022).
@@ 2.2. Spatial Parsing Datasets
Spatial parsing datasets, by contrast, focus on object recognition and relational reasoning within static frames. General-purpose object detection datasets such as COCO (Lin et al., 2014) and Objects365 (Shao et al., 2019) cover a wide spectrum of everyday scenes, while domain-specific datasets—such as TJU-DHD (Pang et al., 2020) for autonomous driving, MVTec AD (Bergmann et al., 2019) and Read-IAD (Wang et al., 2024b) for industrial inspection, and HIOD (Hu et al., 2023) for medical devices—provide high-quality samples that address challenges like occlusion, lighting variations, and scale diversity. Building upon object detection, instance segmentation datasets including ADE20K (Zhou et al., 2017) and COCONut (Deng et al., 2024) offer pixel-level annotations, enabling finer-grained visual understanding.
Beyond object-level tasks, datasets for HOI detection and scene graph generation aim to capture semantic relationships among entities, laying the foundation for high-level visual reasoning (Han et al., 2024). Representative datasets such as HICO-DET (Chao et al., 2018), V-COCO (Gupta and Malik, 2015), Visual Genome (Krishna et al., 2017), and Open Images V4/V6 (Kuznetsova et al., 2020) provide comprehensive annotations of both human-object and object-object interactions. Despite these advances, most existing datasets lack multi-granularity behavior annotations and task-level semantic alignment (Sener et al., 2022), which limits their effectiveness in complex procedural modeling and causal behavior inference (Bi et al., 2021). In contrast, PhysLab supports both temporal and spatial parsing, providing rich annotations that enable a wide range of tasks from fine-grained temporal action segmentation to frame-level relational reasoning.
@@ 3.1. Data Collection
For video acquisition, we employed the Newmine Q40 high-definition camera, which supports 4K resolution (3840×2160), a frame rate of 30 FPS, and a 90-degree distortion-free lens. This device delivers excellent image clarity, wide field-of-view coverage, and stable long-duration performance, making it well-suited for continuous deployment in laboratory settings and autonomous data collection scenarios.
@@ 3.2. Data Annotations
In collaboration with the physics experiment center, we established a professional annotation team and adopted a “cross-annotation & multi-round verification” protocol (Yang et al., 2024) to ensure high annotation quality and consistency. Both temporal and spatial annotations were performed to support multi-granularity visual parsing tasks. The annotated category schema is illustrated in Figure 2, with detailed quantitative statistics provided in our open-source website.
Temporal Annotations. For each experimental video, we annotated the category, start time, and end time of each operational step along the timeline. These annotations facilitate tasks such as action classification in segmented videos and action localization in untrimmed videos (Ding et al., 2023). All labels strictly adhere to the standardized procedures outlined in official experimental manuals, ensuring semantic consistency across the dataset. To enhance annotation objectivity—particularly in delineating precise temporal boundaries—each video was independently labeled by two experienced annotators using the ELAN (Wittenburg et al., 2006) tool. ELAN’s multi-tier timeline structure enables detailed, hierarchical descriptions of complex procedural behaviors. Following independent annotations, results were merged and manually reconciled to address discrepancies and ensure temporal coherence. In total, 3,873 action instances were annotated, each averaging approximately 20 seconds in duration and spanning 32 distinct types of experimental steps. These comprehensive annotations provide a robust foundation for procedural behavior analysis.
Spatial Annotations. To support fine-grained spatial parsing, we extracted keyframes from the experimental videos and eliminated redundant frames using image similarity filtering techniques, resulting in a diverse subset of 4,500 high-quality frames. On these frames, we annotated bounding boxes for 34 categories of experimental instruments and objects, defined by the official equipment specifications in the experimental manuals. In addition, 24 types of interaction verbs were used to describe the functional relationships between the operator and the instruments, based on the semantics of experimental actions. Each annotated frame was thus structured into a triplet of the form ¡Operator, Interaction Verb, Instrument¿. This annotation design enables HOI detection tasks (Yu et al., 2023) and supports the construction of interaction graphs (Zhu et al., 2024) for advanced reasoning and procedural understanding. Given the high density of objects, frequent interactions, and annotation complexity typical of physics experiment scenarios, a dedicated labeling team consisting of 5 graduate students and 18 undergraduate students carried out the annotation process using the Labelme tool. To ensure annotation quality, we implemented multi-round cross-checking and sample-based review procedures (Zhou et al., 2025b).
@@
Table 1. Comparison of procedural visual parsing video datasets. AR: action recognition, OD: object detection, IS: instance segmentation, HOID: human-object interaction detection, Flex.: more than one single execution order for the task, PEs: procedural errors.
Tasks Procedural complexity Dataset Size
Dataset Year Domain/Environment AR OD IS HOID Flex. PEs Videos Hours
@@
3.3. Data Statistics
With the advancement of visual parsing technologies, research datasets have progressively evolved from capturing simple semantic scenes to encompassing more complex and cognitively demanding procedural task videos (Shen and Elhamifar, 2024). Table 1 presents a summary and comparison of representative procedural video datasets, highlighting the advantages and distinctive features of the PhysLab dataset developed in this study.
@@ 3.3. Data Statistics
More recently, research has shifted toward industrial and simulated assembly scenarios, as seen in datasets such as Assembly101 (Sener et al., 2022), HA4M (Cicirelli et al., 2022), ATTACH (Aganian et al., 2023), and IndustReal (Schoonbeek et al., 2024). These datasets feature more intricate task flows, typically involving furniture, toy, or industrial device assembl. They include annotations at both temporal and spatial levels, which significantly enhance their descriptive capacity and support fine-grained analysis. Building upon this trend, the PhysLab dataset we proposed is the first to target the domain of smart education. It comprises 620 real-world physics experiment videos performed by university students, amounting to approximately 31 hours of footage, placing it among the upper-tier procedural video datasets in terms of scale. More importantly, PhysLab introduces the following distinctive features:
@@
In summary, PhysLab stands out for its comprehensive coverage of experimental procedures, behavioral complexity, and fine-grained annotations, making it a high-quality, multi-purpose dataset well suited for downstream applications in experimental process monitoring and visual process analysis.

### v2

@@ Abstract.
Visual parsing of images and videos is critical for a wide range of real-world applications. However, progress in this field is constrained by limitations of existing datasets: (1) limited annotation diversity, which limits the support for diverse vision tasks within a unified dataset; (2) insufficient coverage of domains, particularly a lack of datasets tailored for educational scenarios; and (3) a lack of explicit procedural guidance, with weak logical rules and insufficient representation of a structured task process. To address these gaps, we introduce PhysLab, the first dataset that captures students conducting complex physics experiments. The dataset includes four representative experiments that feature diverse scientific instruments and rich human-object interaction (HOI) patterns. PhysLab comprises 620 long-form videos and provides multi-granularity annotations that support a variety of vision tasks, including action recognition, object detection, HOI analysis, etc. We establish baselines and perform extensive evaluations to highlight key challenges in the parsing of procedural educational videos. We expect PhysLab to serve as a valuable resource for advancing comprehensive visual parsing, facilitating intelligent classroom systems, and fostering closer integration among computer vision, multimedia, and educational technologies. The dataset and the evaluation toolkit are publicly available at https://github.com/ZMH-SDUST/PhysLab.
Procedural Video; Physics Lab Education; Visual Parsing; Multi-Granularity Annotation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145//3746027.3758221ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer vision problems
@@ 1. Introduction
Visual parsing is a fundamental research area in multimedia, encompassing a wide range of tasks such as action recognition, object detection, and HOI detection (Yu et al., 2022). It plays a pivotal role in real-world applications, including surveillance, autonomous driving, and industrial inspection. In recent years, the field has achieved remarkable progress, driven by advances in deep learning and the availability of large-scale annotated datasets (Zhou et al., 2025a). However, this data-centric paradigm also exposes a strong reliance on high-quality and diverse annotations. Although existing datasets are large in scale, they remain short in supporting multi-task collaborative modeling and rich semantic understanding (Li et al., 2024). Specifically, they are often tailored to single-task learning or provide only coarse-grained labels, lacking the structured and detailed annotations needed to jointly model actions, objects, and interactions (Kim et al., 2025). Furthermore, most datasets are predominantly built around common scenarios, such as households (Tang et al., 2019), streets (Liu et al., 2024), and kitchens (Lee et al., 2024), while neglecting cognitively demanding domains like education. As a result, many baseline models tend to prioritize static appearance cues or short-term visual patterns, falling short in modeling more explicit causal relationships, long-term temporal dependencies, and strict physical constraints that are characteristic of complex procedural tasks (Wu et al., 2024). These limitations hinder the development of models capable of deeper and human-like understanding.
Figure 1. An overview of the content of our proposed PhysLab dataset. PhysLab focuses on experimental tasks conducted in complex physics laboratory settings. 1) The top row shows the data collection process and representative data samples. 2) The second row shows the distinctive operational instructions associated with physics experiments, including detailed procedures, instrument usage, and process models in Petri Net format (lower-left), along with multi-granularity annotations (lower-right).
@@
Compared to conventional visual scenes, physics experiments exhibit significantly greater spatiotemporal complexity and cognitive demands (Latif et al., 2024). Experimental procedures typically involve multi-step operations, instrument configurations, and parameter adjustments, all governed by strict task dependencies and procedural order (Zou et al., 2024). These tasks also feature frequent and detailed human-instrument interactions, including tool switching and feedback observation. Such interactions require nuanced recognition and cross-modal reasoning (Huang et al., 2024). Additionally, experimental processes are often underpinned by implicit physical knowledge and causal dynamics (Pols and Dekkers, 2024) that cannot be effectively captured by visual features alone. To facilitate research on such complex and cognitively rich scenarios, we introduce PhysLab, the first multi-granularity annotated video dataset specifically designed for physics experiment instruction.
@@
As shown in Figure 1, PhysLab focuses on four representative undergraduate physics experiments, covering a diverse set of scientific instruments and typical student behaviors. The dataset contains 620 long-form videos (totalling 31 hours), each documenting authentic student experiment sessions. Videos are segmented into action clips, and richly annotated with hierarchical labels including action categories, object locations, interaction relationships, and operation sequences. These annotations provide a unified platform for multi-level or multi-task learning. In addition, the dataset incorporates structured metadata encoding experimental types, procedural steps, and tools used. This facilitates advanced tasks such as procedural modeling and experimental state reasoning. A key concept in this context is the process model (Peterson, 1977), which refers to a structured representation of the experimental workflow, capturing the logical and temporal constraints among operations. It helps in understanding how experiments unfold over time, supporting reasoning about progress, correctness, and deviations during task execution.
@@
To systematically assess the challenges introduced by PhysLab, we conducted extensive benchmark experiments spanning multiple mainstream visual parsing tasks. The evaluation results highlight the dataset’s complexity and expose substantial limitations in current methods, particularly in terms of insufficient visual context fusion, lack of prior knowledge, and pronounced inter-class performance disparity. We believe PhysLab offers a valuable foundation for advancing research in complex behavior modeling, improving educational process analysis, and bridging the gap between visual understanding and intelligent educational systems.
@@ 2.1. Temporal Parsing Datasets
Temporal parsing datasets focus on modeling the temporal dynamics of videos and support tasks such as action classification, temporal action localization, and action segmentation (Sun et al., 2022; Wang et al., 2023). Early datasets such as UCF101 (Soomro et al., 2012), HMDB51 (Kuehne et al., 2011), and Kinetics (Carreira and Zisserman, 2017) contain a variety of short and edited clips depicting everyday human actions, and have been widely adopted for action classification tasks. However, these datasets offer limited diversity in action types and lack well-defined temporal boundaries. To address this, subsequent datasets like THUMOS14 (Idrees et al., 2017), ActivityNet1.3 (Caba Heilbron et al., 2015), and FineAction (Liu et al., 2022) introduced untrimmed videos with denser temporal annotations, enabling research on action detection and segmentation.
More recently, as the community shifts towards modeling structured activities, several procedural video datasets have emerged. Datasets such as Breakfast (Kuehne et al., 2014), COIN (Tang et al., 2019), and Assembly101 (Sener et al., 2022) focus on multi-step instructional tasks with well-defined goals, supporting the study of procedural understanding. In addition, datasets like EgoPER (Lee et al., 2024), CaptainCook4D (Peddi et al., 2024), and IndustReal (Schoonbeek et al., 2024) incorporate annotations of execution errors and deviations, facilitating research on anomaly detection in task workflows. Despite recent progress, existing datasets lack explicit procedural guidance for modeling complex workflows and are largely confined to narrow domains such as cooking and mechanical assembly (Tan et al., 2021). Datasets tailored to student laboratory experiments remain scarce, yet they are essential for advancing research at the intersection of visual perception and intelligent educational systems (Piloto et al., 2022).
@@ 2.2. Spatial Parsing Datasets
Spatial parsing datasets focus on object recognition and relational reasoning within static frames. General object detection datasets, such as COCO (Lin et al., 2014) and Objects365 (Shao et al., 2019), cover daily scenes. Domain-specific datasets, such as TJU-DHD (Pang et al., 2020) for autonomous driving, MVTec AD (Bergmann et al., 2019) and Read-IAD (Wang et al., 2024b) for industrial inspection, and HIOD (Hu et al., 2023) for medical devices, provide high-quality samples that address challenges like occlusion, lighting variations, and scale diversity. Building upon object detection, instance segmentation datasets including ADE20K (Zhou et al., 2017) and COCONut (Deng et al., 2024) offer pixel-level annotations, enabling finer-grained visual understanding.
Beyond object recognition tasks, datasets for HOI detection and scene graph generation aim to capture semantic relationships among objects, enabling more comprehensive contextual and relational reasoning (Han et al., 2024). Representative datasets such as HICO-DET (Chao et al., 2018), V-COCO (Gupta and Malik, 2015), Visual Genome (Krishna et al., 2017), and Open Images V4/V6 (Kuznetsova et al., 2020) provide comprehensive annotations of both human-object and object-object interactions. Despite these advances, most existing datasets lack multi-granularity behavior annotations and task-level semantic alignment (Sener et al., 2022), which limits their effectiveness in complex procedural modeling and causal behavior inference (Bi et al., 2021). In contrast, PhysLab integrates comprehensive temporal and spatial annotations, enabling diverse tasks ranging from video-level temporal action segmentation to frame-level object relational reasoning.
@@ 3.1. Data Collection
For video acquisition, we used the Newmine Q40 high-definition camera, which supports 4K resolution (3840×2160), a frame rate of 30 FPS, and a 90-degree distortion-free lens. This device delivers excellent image clarity, wide field-of-view coverage, and stable long-duration performance, making it well-suited for continuous laboratory deployment and autonomous data collection.
@@ 3.2. Data Annotations
In collaboration with the physics experiment center, we established a professional annotation team and adopted a “multi-annotator & multi-round verification” protocol (Yang et al., 2024) to ensure high annotation quality and consistency. Each sample was first annotated by an assigned annotator, then iteratively refined through reviews by multiple annotators. Both temporal and spatial annotations were performed to support multi-granularity visual parsing tasks. The annotated category schema is illustrated in Figure 2, with detailed quantitative statistics provided on our open-source website.
@@
Temporal Annotations. For each video, we annotated the category, start time, and end time of each operational step along the timeline. These annotations facilitate tasks such as action classification in segmented videos and action localization in untrimmed videos (Ding et al., 2023). All labels strictly adhere to standardized procedures outlined in official experimental manuals, ensuring semantic consistency across the dataset. To enhance annotation objectivity, particularly in delineating precise temporal boundaries, each video was independently labeled by two experienced annotators using the ELAN (Wittenburg et al., 2006) tool. ELAN’s multi-tier timeline structure enables detailed, hierarchical descriptions of complex procedural behaviors. Following independent annotations, results were merged and manually reconciled to address discrepancies and ensure temporal coherence. In total, 3,873 action instances were annotated, each averaging approximately 20 seconds in duration and spanning 32 distinct types of experimental steps. These comprehensive annotations provide a robust foundation for procedural behavior analysis.
Spatial Annotations. To support spatial parsing, we extracted keyframes from the experimental videos and eliminated redundant frames using image similarity filtering techniques, resulting in a diverse subset of 4,500 high-quality frames. On these frames, we annotated bounding boxes for 34 categories of experimental instruments and objects, defined by the official equipment specifications in the experimental manuals. In addition, 24 types of interaction verbs were used to describe the functional relationships between the operator and the instruments, based on the semantics of experimental actions. Each annotated frame was thus structured into a triplet of the form ¡Operator, Interaction Verb, Instrument¿. This annotation design enables HOI detection tasks (Yu et al., 2023) and supports the construction of interaction graphs (Zhu et al., 2024) for advanced reasoning and procedural understanding. Given the high density of objects, frequent interactions, and annotation complexity typical of physics experiment scenarios, a dedicated labeling team consisting of 5 postgraduate students and 18 undergraduate students carried out the annotation process using the Labelme tool. To ensure annotation quality, we implemented multi-round cross-checking and sample-based review procedures (Zhou et al., 2025b).
3.3. Data Statistics
With the advancement of visual parsing technologies, research datasets have progressively evolved from capturing simple semantic scenes to encompassing more complex and cognitively demanding procedural task videos (Shen and Elhamifar, 2024). Table 1 presents a summary and comparison of representative procedural video datasets, highlighting the advantages and distinctive features of the PhysLab dataset.
Table 1. Comparison of procedural visual parsing video datasets. AC: action classification, TAPG: temporal action proposal generation, AL: action alignment, AS: action segmentation, OD: object detection, IS: instance segmentation, OR: occlusion restoration, HOID: HOI detection, Flex.: more than one single execution order for the task, PEs: procedural errors.
Tasks Complexity Dataset Size
Dataset Year Domain/Environment AC TAPG AL AS OD IS OR HOID Flex. PEs Videos Hours
@@ 3.3. Data Statistics
More recently, research has shifted toward industrial and simulated assembly scenarios, as seen in datasets such as Assembly101 (Sener et al., 2022), HA4M (Cicirelli et al., 2022), ATTACH (Aganian et al., 2023), and IndustReal (Schoonbeek et al., 2024). These datasets feature more intricate task flows, typically involving furniture, toy, or industrial device assembly. They include annotations at both temporal and spatial levels, which significantly enhance their descriptive capacity and support comprehensive visual analysis. Building upon this trend, the PhysLab dataset we proposed is the first to target the domain of smart education. It comprises 620 real-world physics experiment videos performed by university students, totaling approximately 31 hours of footage, which is a considerable duration for datasets in this field. More importantly, PhysLab introduces the following distinctive features:
@@
In summary, PhysLab stands out for its comprehensive coverage of experimental procedures, behavioral complexity, and multi-granularity annotations, making it a high-quality and multi-purpose dataset well suited for downstream applications in experimental process monitoring and visual process analysis.